OpenStudy (anonymous):

Please Help!! If x = 7 sin θ, write the expression θ/2 − sin 2θ/4 in terms of just x. (Simplify the double angle using a double-angle identity.)

OpenStudy (freckles):

did you make sure you put all the ( ) in the right place?

OpenStudy (freckles):

see my post on your previous question

OpenStudy (anonymous):

I did! I did! it was wrong....

OpenStudy (anonymous):

i got it we did the whole thing wrong

OpenStudy (freckles):

I don't see how we did

OpenStudy (freckles):

unless again you didn't mean \[\frac{\theta}{2}-\frac{\sin(2 \theta)}{4}\]

OpenStudy (freckles):

did you mean \[\frac{\theta}{2}-\sin(\frac{2 \theta}{4})?\]

OpenStudy (anonymous):

yes but the way i inserted the info into this problem was wrong